# references.bib

### v1

@@ 4 Properties of $3$ variable quasi-invariants
The proof is largely the same as in [wang2022explicit]; the only difference is in the last step. Namely, now we have 2 2-dimensional indecomposable representations $\mathrm{sign-triv}$ and $\mathrm{triv-sign},$ but an element in the $(-1)$-eigenspace of $s_{12}$ in $\mathrm{triv-sign}$ must be in a copy of $\mathrm{sign}$. ∎
@@
Let $K$ be a generator of $Q_{m}(3,\mathbf{F}_{3})$ and write $K=(x_{i}-x_{j})^{2m+1}K^{\prime}$. Then $K^{\prime}$ is not divisible by any nonconstant symmetric polynomial.
The proof of this corollary is identical to the proof of Corollary 3.3 in [wang2022explicit].
@@
Let $K$ and $L$ be distinct generators of $Q_{m}(3,\mathbbm{k})_{\mathrm{sign-triv}},$ and let $V$ and $W$ be the copies of $\mathrm{sign-triv}$ generated by $K$ and $L$ respectively. Then $Ks_{23}L+Ls_{23}K$ is a nonzero element of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ and $\deg V+\deg W\geq 6m+3$.
Noting that $\wedge^{2}(\mathrm{sign-triv})=\mathrm{sign}$, the proof of this lemma is also identical to the proof of Lemma 3.4 in [wang2022explicit].
Lemma 3.5 from [wang2022explicit] does not completely hold in characteristic $3$. A very similar and useful version does, however:
@@
We note that $(L+s_{23}L)K-(K+s_{23}K)L=c\prod_{i<j}(x_{i}-x_{j})^{2m+1}$ for some $c\neq 0$ by Lemma 4.3. Moreover, $L+s_{23}L$ and $K+s_{23}K$ are symmetric because $K$ and $L$ are both acted on by negation by $s_{12},$ so elements in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ are generated by $K$ and $L.$ From there, the fact that $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by $K,$ $L,$ and $1$ over $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ follows from the first part of the proof from [wang2022explicit].
To prove freeness, assume for the sake of contradiction that there was a relation $PK+QL+S=0$ for symmetric polynomials $P,$ $Q,$ and $S.$ $PK$ and $QL$ are both in the $(-1)$-eigenspace of $s_{12}$ while $S$ is not, so $S=0.$ Thus we have $PK=-QL$ and from there we can proceed the same as [wang2022explicit].
@@ 5 Ren–Xu counterexamples
In [Ren_2020], Ren and Xu found polynomials of the form $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ in $Q_{m}(3,\mathbf{F}_{3})$ with degree less than $13m+$ 1 where $P_{k}$ is the map of the $13k+$ 1 degree generator of $Q_{k}(3,\mathbf{Q})$ into characteristic $3$ and where $a,$ $k,$ and $b$ are natural numbers. We refer to these polynomials as Ren–Xu counterexamples as they demonstrate the Hilbert series of $Q_{m}(3,\mathbf{F}_{3})$ differs from that of $Q_{m}(3,\mathbf{Q})$ for certain $m.$
@@
A key step in describing the Hilbert series of $Q_{m}(3,\mathbf{F}_{3})$ is proving [Ren_2020]’s conjecture:
Conjecture 5.2 ([Ren_2020]).
If the Hilbert series of $Q_{m}(3,\mathbf{F}_{p})$ differs from that of $Q_{m}(3,\mathbf{Q}),$ then there exists integers $a\geq 0$ and $k\geq 0$ such that
@@
The main step for proving the conjecture is the following theorem.
@@
$Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is either freely generated by a generator of degree $3m+1,$ $23m+$ 2, and the polynomial $1$, or it is freely generated by $R_{m},$ another generator in degree $6m+3-\deg R_{m},$ and the polynomial $1$.
@@
If $m\in X,$ we must have $R_{m}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $a,b,k$ are natural numbers and $k\not\in X$.
@@
Assume for contradiction that there exists a nonnegative integer $m\in X$ such that $R_{m}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $a,b,k$ are natural numbers and $k\in X$. Then if $R_{k}=P_{l}^{3^{c}}\prod(x_{i}-x_{j})^{2d},$ the polynomial
${R_{k}}^{3^{a}}\prod(x_{i}-x_{j})^{2b}=P_{l}^{3^{a+c}}\prod(x_{i}-x_{j})^{2d\cdot 3^{a}+2b}$
@@
This lemma allows us to consider only counterexamples $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ such that $Q_{k}(3,\mathbf{F}_{3})$ does not contain a Ren–Xu counterexample.
From [Ren_2020], the Hilbert series for $Q_{m}(3,\mathbf{F}_{3})$ differs from characteristic 00 when there exists $a\in\mathbf{N}_{0}$ such that
$\frac{1}{3}\leq\left\{\frac{m}{3^{a}}\right\}\leq\frac{1}{3}-\frac{1}{3^{a}}.$
Notice this is equivalent to $m\pmod{3^{a}}$ being in $\{3^{a-1},3^{a-1}+1,\dotsm,2\cdot 3^{a-1}-1\}$.
@@
Suppose $m$ had the digit $1$ in the $a$th position from the right. Then $m\pmod{3^{a}}$ has a leading digit of $1$ if we choose $m\pmod{3^{a}}$ to be between 00 and $13^{a}-$ 1 inclusive. However, this implies that $m$ is in $\{3^{a-1},3^{a-1}+1,\dotsm,2\cdot 3^{a-1}-1\}$, so $m$ is a counterexample. ∎
@@
$m=\sum_{i=0}c_{i}3^{i}$
where $c_{i}$ is 00 or $2$. Thus $m$ must be even. ∎
@@
By Corollary 5.6, if $m\not\in X,$ $m$ is even. Then $m+1$ is odd, so by the contrapositive of Corollary 5.6, $m+1\in X.$
@@
If $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $3m+1,$ then $m+1\in X$ and $\deg R_{m+1}=3m+3$.
@@
Because $\deg R_{m+1}<3m+4$ and $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}\subset Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, we have $43m+1\leq\deg R_{m+1}<3m+$ 4 . By construction $3|\deg R_{m+1},$ so $\deg R_{m+1}=3m+3.$
@@
Suppose $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a smallest degree generator $L$ in degree $13m+$ 1 . Assume that for all $i<m$, if $i\not\in X$, then $Q_{i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a degree $13i+$ 1 generator. Then $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has no nonsymmetric degree $13m+$ 1 or $23m+$ 2 element.
@@
Any nonsymmetric $13m+$ 1 degree element in $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ must be a scalar multiple of $L,$ so assume for contradiction $L$ is in $Q_{m+1}(3,\mathbf{F}_{3}).$ By Lemma 5.8, $R_{m+1}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ is in degree $33m+$ 3 for natural numbers $a,$ $b,$ $k.$ By Lemma 5.4, $k\not\in X$ implying $P_{k}$ is a $13k+$ 1 generator of $Q_{k}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ using our assumption. Moreover, with any other generator in a degree less than $33m+$ 3 violating Lemma 4.3, $R_{m+1}$ must be generated by $L,$ so $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}=SL$ for some degree $2$ symmetric polynomial $S$. A degree $2$ symmetric polynomial divisible by $(x_{i}-x_{j})$ is impossible, so $S|P_{k}^{3^{a}}$ which implies either $S|P_{k}$ or $(x_{1}+x_{2}+x_{3})|P_{k}.$ If we let $P_{k}=P_{k}^{\prime}(x_{1}-x_{2})^{2k+1},$ in both cases $S|P_{k}^{\prime}$ or $(x_{1}+x_{2}+x_{3})|P_{k}^{\prime}$. However, by our assumption $P_{k}$ is a generator, so $P_{k}^{\prime}$ is not divisible by any nonconstant symmetric polynomial by Corollary 4.2.
Similarly, suppose for contradiction that $K$ is a nonsymmetric element of $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ of degree $23m+$ 2. Since $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has no nonsymmetric $13m+$ 1 degree element, $K$ must be a generator. By Lemma 4.3, $K$ is the only generator in degree less than $3m+3,$ so $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ is symmetric polynomial multiple of $K.$ However the only symmetric polynomials of degree $1$ are multiples of $x_{1}+x_{2}+x_{3},$ implying $(x_{1}+x_{2}+x_{3})|P_{k}$ which is impossible by Corollary 4.2. ∎
Note that by [felder2001action], $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ has generators in degree $13m+$ 1 and $3m+2,$ and by [wang2022explicit], such generators with even degree are divisible by $x_{1}+x_{2}-2x_{3}.$ Let $\pi$ be the canonical mapping from characteristic 00 to characteristic $3$. We then have the following lemma:
@@
$G=(x_{1}+x_{2}+x_{3})\left(\frac{K^{\prime}-L^{\prime}}{3}\right)-x_{3}K^{\prime},$
@@
$\pi(G)=(x_{1}+x_{2}+x_{3})\pi\left(\frac{K^{\prime}-L^{\prime}}{3}\right)-x_{3}L$
@@
A higher degree generator of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ has degree $23m+$ 2 . With $\deg L=3m+1$ implying $m\not\in X$, $23m+$ 2 is even by Corollary 5.6. Using [wang2022explicit], we let $(x_{1}+x_{2}-2x_{3})K^{\prime}$ be an arbitrary degree $23m+$ 2 generator for $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ with coprime integer coefficients. Similarly $\pi((x_{1}+x_{2}-2x_{3})K^{\prime})=(x_{1}+x_{2}+x_{3})\pi(K^{\prime})$ is an element of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}},$ so $\pi(K^{\prime})$ is a non-symmetric polynomial of degree $13m+$ 1 in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. Thus it must be a scalar multiple of $L,$ and we may set $\pi(K^{\prime})=L.$
Let $G=(x_{1}+x_{2}+x_{3})\left(\frac{K^{\prime}-L^{\prime}}{3}\right)-x_{3}K^{\prime}.$ Since
$(x_{1}+x_{2}-2x_{3})K^{\prime}-(x_{1}+x_{2}+x_{3})L^{\prime}=(x_{1}+x_{2}+x_{3})(K^{\prime}-L^{\prime})-3x_{3}K^{\prime}$
@@
$\pi(G)=(x_{1}+x_{2}+x_{3})\pi\left(\frac{K^{\prime}-L^{\prime}}{3}\right)-x_{3}L.$
If $\pi(G)$ generated by $L$, we must have $\pi(G)=c(x_{1}+x_{2}+x_{3})L$ for some $c\in\mathbf{F}_{3}$ since $\deg\left(\pi(G)\right)=\deg(L)+1.$ However $x_{1}+x_{2}+x_{3}$ does not divide $x_{3}L$ since $L$ is a generator, so $x_{1}+x_{2}+x_{3}\nmid\pi(G).$ Then if $\pi(G)$ was not a generator, there must be some generator other than $L$ for $Q_{m}(3,\mathbf{F}_{3})$ in degree less than $23m+$ 2 which violates Lemma 4.3. Thus, $\pi(G)$ is a generator. ∎
We aim to prove that minimum Ren–Xu counterexamples are generators and represent the only cases where the Hilbert series of the quasi-invariants differs between characteristics 00 and $3$. To this end, we describe the degree of Ren–Xu counterexamples.
@@
$(x_{1}-x_{2})^{9}$
$(x_{1}-x_{2})^{9}$
$(x_{1}-x_{2})^{9}\prod(x_{i}-x_{j})^{2}.$
We note that since $(x_{1}-x_{2})^{9}\in Q_{4}(3,\mathbf{F}_{3}),$ $(x_{1}-x_{2})^{9}$ is the Ren–Xu counterexample for both $m=3$ and $m=4.$ Moreover, the counterexample in $Q_{5}(3,\mathbf{F}_{3})$ is the previous counterexample $(x_{1}-x_{2})^{9}$ muliplied by $\prod(x_{i}-x_{j})^{2}$ to add the extra factor of $(x_{1}-x_{2})^{2}.$ In this way the degree of counterexample stays constant for the first half of the “staircase” and climbs by $6$ per each increase in $m$ thereafter. Moreover, we note that $m=2,6\not\in X,$ so our “staircase” is surrounded by non-counterexamples. One can also compute another generator for $m=3,$ $4,$ $5$ in degree $12,$ $18,$ and $18$ respectively. Since $9+12=6\cdot 3+3,$ $9+18=6\cdot 4+3,$ and $15+18=6\cdot 5+3,$ $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by each of these generators and $1$ by Lemma 4.4. This way we see that the upper degree generators form a complement to the lower degree ones, climbing by $6$ degrees initially and staying constant for the second half of the staircase.
@@
0010101010202020203030303040404040505050500050505050100100100100150150150150mDegree of generatorGenerator $1$Generator $2$
Figure 1: Degrees of generators in characteristic $3$ with respect to $m$
We prove that Ren–Xu counterexamples follow this staircase pattern:
@@
Let $m$ be a natural number not in $X$ and let $d$ be the largest integer such that $R_{m+1}$ lies in $Q_{m+d}(3,\mathbf{F}_{3})$. Suppose that for all $k\leq m,$ if $k\not\in X$, then $Q_{k}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $13k+$ 1 . Then $R_{m+i}=R_{m+1}$ in degree $33m+$ 3 for $1\leq i\leq d$ and $R_{m+i}=R_{m+1}\prod(x_{i}-x_{j})^{2(i-d)}$ in degree $3m+3+6(i-d)$ for $d<i<2d.$
@@
Let $R_{m+1}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $k\in\mathbb{Z}_{\geq 0},a\in\mathbb{Z}_{>0}$, and $b=\max\left\{\frac{2m+3-3^{a}(2k+1)}{2},0\right\}.$ If $b$ is positive the polynomial $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2(b-1)}$ has degree less than $23m-$ 2 and is at least $m$-quasi-invariant since $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ has degree less than $43m+$ 4 . Thus $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2(b-1)}$ is a Ren–Xu counterexample for $Q_{m}(3,\mathbf{F}_{3}),$ a contradiction.

### v2

@@ 4 Properties of $\boldsymbol{3}$ variable quasi-invariants
The proof is largely the same as in [13]; the only difference is in the last step. Namely, now we have 2 2-dimensional indecomposable representations $\mathrm{sign-triv}$ and $\mathrm{triv-sign}$, but an element in the $(-1)$-eigenspace of $s_{12}$ in $\mathrm{triv-sign}$ must be in a copy of $\mathrm{sign}$. ∎
@@
Let $K$ be a generator of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ in $V_{i_{1}i_{2}}^{-}$ for some copy $V$ of $\mathrm{sign-triv}$ and write $K=(x_{i_{1}}-x_{i_{2}})^{2m+1}K^{\prime}$. Then $K^{\prime}$ is not divisible by any nonconstant symmetric polynomial.
The proof of this corollary is identical to the proof of [13, Corollary 3.3].
@@
Let $K$ and $L$ be distinct generators of $Q_{m}(3,\mathbbm{k})_{\mathrm{sign-triv}}$, and let $V$ and $W$ be the copies of $\mathrm{sign-triv}$ generated by $K$ and $L$ respectively such that $K\in V_{i_{1}i_{2}}^{-}$ and $L\in W_{i_{1}i_{2}}^{-}$. Then $Ks_{23}L-Ls_{23}K$ is a nonzero element of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ and $\deg V+\deg W\geq 6m+3$.
Noting that $\wedge^{2}(\mathrm{sign-triv})=\mathrm{sign}$, the proof of this lemma is also identical to the proof of [13, Lemma 3.4].
Lemma 3.5 from [13] does not completely hold in characteristic $3$. A very similar and useful version does, however, and we have the following.
@@
We note that $(L+s_{23}L)K-(K+s_{23}K)L=c\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m+1}$ for some $c\neq 0$ by Lemma 4.3. Moreover, $L+s_{23}L$ and $K+s_{23}K$ are symmetric because $K$ and $L$ are both acted on by negation by $s_{12}$, so elements in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ are generated by $K$ and $L$. From there, the fact that $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by $K$, $L$, and $1$ over $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ follows from the first part of the proof from [13].
To prove freeness, assume for the sake of contradiction that there was a relation $PK+QL+S=0$ for symmetric polynomials $P$, $Q$, and $S$. $PK$ and $QL$ are both in the $(-1)$-eigenspace of $s_{12}$ while $S$ is not, so $S=0$. Thus we have $PK=-QL$ and from there we can proceed the same as [13].
@@ 5 Ren–Xu counterexamples
In [11], Ren and Xu found polynomials of the form $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$ in $Q_{m}(3,\mathbf{F}_{3})$ with degree strictly less than $13m+$ 1, where $P_{k}$ is the map of the $13k+$ 1 degree generator of $Q_{k}(3,\mathbf{Q})$ into characteristic $3$ and where $a$, $k$, and $b$ are natural numbers. We refer to these polynomials as Ren–Xu counterexamples as they demonstrate the Hilbert series of $Q_{m}(3,\mathbf{F}_{3})$ differs from that of $Q_{m}(3,\mathbf{Q})$ for certain $m$.
@@
A key step in describing the Hilbert series of $Q_{m}(3,\mathbf{F}_{3})$ is proving Ren–Xu’s conjecture [11] for ${n=3}$ and ${p=3}$.
Conjecture 5.2 ([11]).
If the Hilbert series of $Q_{m}(n,\mathbf{F}_{p})$ differs from that of $Q_{m}(n,\mathbf{Q})$, then there exists integers $a\geq 0$ and $k\geq 0$ such that
@@
The main step for proving the conjecture for $n=3$, $p=3$ is the following theorem.
@@
$Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is either freely generated by a generator of degree $13m+$ 1, $23m+$ 2, and the polynomial $1$, or it is freely generated by $R_{m}$, another generator in degree $6m+3-\deg R_{m}$, and the polynomial $1$.
@@
If $m\in X$, we must have $R_{m}=P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$, where $a$, $b$, $k$ are natural numbers and $k\not\in X$.
@@
Assume for contradiction that there exists a nonnegative integer $m\in X$ such that $R_{m}=P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$, where $a$, $b$, $k$ are natural numbers and $k\in X$. Then if $R_{k}=P_{l}^{3^{c}}\prod(x_{i_{1}}-x_{i_{2}})^{2d}$, the polynomial
${R_{k}}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}=P_{l}^{3^{a+c}}\prod(x_{i_{1}}-x_{i_{2}})^{2d\cdot 3^{a}+2b}$
@@
This lemma allows us to consider only counterexamples $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$ such that $Q_{k}(3,\mathbf{F}_{3})$ does not contain a Ren–Xu counterexample.
From [11], the Hilbert series for $Q_{m}(3,\mathbf{F}_{3})$ differs from characteristic 00 when there exists $a\in\mathbf{N}_{0}$ such that
$\frac{1}{3}\leq\Bigl{\{}\frac{m}{3^{a}}\Bigr{\}}\leq\frac{2}{3}-\frac{1}{3^{a}}.$
Notice this is equivalent to $m\pmod{3^{a}}$ being in $\bigl{\{}3^{a-1},3^{a-1}+1,\dots,2\cdot 3^{a-1}-1\bigr{\}}$.
@@
Suppose $m$ had the digit $1$ in the $a$-th position from the right. Then $m\pmod{3^{a}}$ has a leading digit of $1$ if we choose $m\pmod{3^{a}}$ to be between 00 and $13^{a}-$ 1 inclusive. However, this implies that $m\pmod{3^{a}}$ is in $\bigl{\{}3^{a-1},3^{a-1}+1,\dots,2\cdot 3^{a-1}-1\bigr{\}}$, so $m$ is a counterexample. ∎
@@
$m=\sum_{j=0}c_{j}3^{j},$
where $c_{j}$ is 00 or $2$. Thus $m$ must be even. ∎
@@
By Corollary 5.6, if $m\not\in X$, $m$ is even. Then $m+1$ is odd, so by the contrapositive of Corollary 5.6, $m+1\in X$. ∎
@@
If $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $13m+$ 1, then $m+1\in X$ and $\deg R_{m+1}=3m+3$.
@@
Because $\deg R_{m+1}<3m+4$ and $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}\subset Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, we have $43m+1\leq\deg R_{m+1}<3m+$ 4. By construction $3|\deg R_{m+1}$, so $\deg R_{m+1}=3m+3$. ∎
@@
Suppose $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a smallest degree generator $L$ in degree $13m+$ 1. Assume that for all $j<m$, if $j\not\in X$, then $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a degree $13j+$ 1 generator. Then $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has no nonsymmetric degree $13m+$ 1 or $23m+$ 2 element.
@@
Any nonsymmetric $13m+$ 1 degree element in $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ must be a scalar multiple of $L$, so assume for contradiction $L$ is in $Q_{m+1}(3,\mathbf{F}_{3})$. By Lemma 5.8, $R_{m+1}=P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$ is in degree $33m+$ 3 for natural numbers $a$, $b$, $k$. By Lemma 5.4, $k\not\in X$ implying $P_{k}$ is a $13k+$ 1 generator of $Q_{k}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ using our assumption. Moreover, with any other generator in a degree less than $33m+$ 3 violating Lemma 4.3, $R_{m+1}$ must be generated by $L$, so $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}=SL$ for some degree $2$ symmetric polynomial $S$. A degree $2$ symmetric polynomial divisible by $(x_{i_{1}}-x_{i_{2}})$ is impossible, so $S|P_{k}^{3^{a}}$ which implies either $S|P_{k}$ or $(x_{1}+x_{2}+x_{3})|P_{k}$. Since $\overline{P_{k}}$ is in the $(-1)$-eigenspace of $s_{12}$, $P_{k}$ is as well and by Lemma 4.1 we have $P_{k}=P_{k}^{\prime}(x_{1}-x_{2})^{2k+1}$. In both cases either $S|P_{k}^{\prime}$ or $(x_{1}+x_{2}+x_{3})|P_{k}^{\prime}$. However, by our assumption $P_{k}$ is a generator, so $P_{k}^{\prime}$ is not divisible by any nonconstant symmetric polynomial by Corollary 4.2.
Similarly, suppose for contradiction that $K$ is a nonsymmetric element of $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ of degree $23m+$ 2. Since $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has no nonsymmetric $13m+$ 1 degree element, $K$ must be a generator. By Lemma 4.3, $K$ is the only generator in degree less than $33m+$ 3, so $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$ is a symmetric polynomial multiple of $K$. However, the only symmetric polynomials of degree $1$ are multiples of $x_{1}+x_{2}+x_{3}$, implying $(x_{1}+x_{2}+x_{3})|P_{k}$ which is impossible by Corollary 4.2. ∎
Note that by [9], $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ has generators in degree $13m+$ 1 and $23m+$ 2, and by [13], such generators with even degree are divisible by $x_{1}+x_{2}-2x_{3}$. Let $\pi$ be the canonical mapping from characteristic 00 to characteristic $3$. We then have the following lemma.
@@
$G=(x_{1}+x_{2}+x_{3})\biggl{(}\frac{K^{\prime}-L^{\prime}}{3}\biggr{)}-x_{3}K^{\prime},$
@@
$\pi(G)=(x_{1}+x_{2}+x_{3})\pi\biggl{(}\frac{K^{\prime}-L^{\prime}}{3}\biggr{)}-x_{3}L$
@@
A higher degree generator of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ has degree $23m+$ 2. With $\deg L=3m+1$ implying $m\not\in X$, $23m+$ 2 is even by Corollary 5.6. Using [13], we let $(x_{1}+x_{2}-2x_{3})K^{\prime}$ be an arbitrary degree $23m+$ 2 generator for $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ with coprime integer coefficients. Similarly, $\pi((x_{1}+x_{2}-2x_{3})K^{\prime})=(x_{1}+x_{2}+x_{3})\pi(K^{\prime})$ is an element of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, so $\pi(K^{\prime})$ is a non-symmetric polynomial of degree $13m+$ 1 in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. Thus it must be a scalar multiple of $L$, and we may set $\pi(K^{\prime})=L$.
Let $G=(x_{1}+x_{2}+x_{3})\big{(}\frac{K^{\prime}-L^{\prime}}{3}\big{)}-x_{3}K^{\prime}$. Since
$(x_{1}+x_{2}-2x_{3})K^{\prime}-(x_{1}+x_{2}+x_{3})L^{\prime}=(x_{1}+x_{2}+x_{3})\big{(}K^{\prime}-L^{\prime}\big{)}-3x_{3}K^{\prime}$
@@
$\pi(G)=(x_{1}+x_{2}+x_{3})\pi\biggl{(}\frac{K^{\prime}-L^{\prime}}{3}\biggr{)}-x_{3}L.$
If $\pi(G)$ generated by $L$, we must have $\pi(G)=c(x_{1}+x_{2}+x_{3})L$ for some $c\in\mathbf{F}_{3}$ since $\deg\left(\pi(G)\right)=\deg(L)+1$. However, $x_{1}+x_{2}+x_{3}$ does not divide $x_{3}L$ since $L$ is a generator, so $x_{1}+x_{2}+x_{3}\nmid\pi(G)$. Then if $\pi(G)$ was not a generator, there must be some generator other than $L$ for $Q_{m}(3,\mathbf{F}_{3})$ in degree less than $23m+$ 2 which violates Lemma 4.3. Thus, $\pi(G)$ is a generator. ∎
We aim to prove that minimum Ren–Xu counterexamples are generators and represent the only cases, where the Hilbert series of the quasi-invariants differs between characteristics 00 and $3$. To this end, we describe the degree of Ren–Xu counterexamples.
@@
$(x_{1}-x_{2})^{9},\qquad(x_{1}-x_{2})^{9},\qquad(x_{1}-x_{2})^{9}\prod(x_{i_{1}}-x_{i_{2}})^{2}.$
We note that since $(x_{1}-x_{2})^{9}\in Q_{4}(3,\mathbf{F}_{3})$, $(x_{1}-x_{2})^{9}$ is the Ren–Xu counterexample for both ${m=3}$ and $m=4$. Moreover, the counterexample in $Q_{5}(3,\mathbf{F}_{3})$ is the previous counterexample $(x_{1}-x_{2})^{9}$ multiplied by $\prod(x_{i_{1}}-x_{i_{2}})^{2}$ to add the extra factor of $(x_{1}-x_{2})^{2}$. In this way the degree of counterexample stays constant for the first half of the “staircase” and climbs by $6$ per each increase in $m$ thereafter. Moreover, we note that $m=2,6\not\in X$, so our “staircase” is surrounded by non-counterexamples. One can also compute another generator for $m=3,4,5$ in degree $12$, $18$, and $18$ respectively. Since $9+12=6\cdot 3+3$, $9+18=6\cdot 4+3$, and $15+18=6\cdot 5+3$, $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by each of these generators and $1$ by Lemma 4.4. This way we see that the upper degree generators form a complement to the lower degree ones, climbing by $6$ degrees initially and staying constant for the second half of the staircase.
@@
[…]
Figure 1: Degrees of generators in characteristic $3$ with respect to $m$.
We prove that Ren–Xu counterexamples follow this staircase pattern.
@@
Let $m$ be a natural number not in $X$ and let $d$ be the largest integer such that $R_{m+1}$ lies in $Q_{m+d}(3,\mathbf{F}_{3})$. Suppose that for all $k\leq m$, if $k\not\in X$, then $Q_{k}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $13k+$ 1. Then $R_{m+j}=R_{m+1}$ in degree $33m+$ 3 for $1\leq j\leq d$ and $R_{m+j}=R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(j-d)}$ in degree $3m+3+6(j-d)$ for $d<j<2d$.
@@
where $k$ is a nonnegative integer, $a$ is a positive integer, and $b=\max\bigl{\{}0,\frac{2m+3-3^{a}(2k+1)}{2}\bigr{\}}$. If $b$ is positive, the polynomial $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2(b-1)}$ has degree less than $23m-$ 2 and is at least $m$-quasi-invariant since $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$ has degree less than $43m+$ 4. Thus $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2(b-1)}$ is a Ren–Xu counterexample for $Q_{m}(3,\mathbf{F}_{3})$, a contradiction.